On September 30, 20x6, Riches Inc. made three investments: 1369 shares of Wendle Inc., 4140 shares of Sam International, and 807 shares of Hinter Inc. The shares of Wendle Inc. and Hinter Inc. are classified as FVTPL and the shares of Sam International are 0o classified as FVTOCI. The market price per share is as follows: Wendle Inc Hinter Inc Sam International S5.25 $8.47 Sept. 30, x6 $13.00 Dec. 31, x6 $16.39 $7.36 $4.82 What amount would be reported as total comprehensive income for the year 20x6, assuming income before gains/losses on investments is $1649721 Select one: O a. S167563 b. $180894 c.$15922 d. $13331 Check
Answer #1

Amount of unrealised gain or (loss) on shares of Wendle Inc. = (16.39 - 13) * 1369

= 3.39 * 1369

= $4,641

Amount of unrealised gain or (loss) on shares of Sam International = (8.47 - 5.25) * 4140

= 3.22 * 4140

= $13,331

Amount of unrealised gain or (loss) on shares of Hinter Inc. = (4.82 - 7.36) * 807

= -2.54 * 807

= $(2,050)

Amount to be included in Profit and Loss Account = $4,641 + $(2,050)

= $2,592

Amount to be included in Other Comprehensive Income Account = $13,331

Total comprehensive income is the combination of Profit and Loss Account and Other Comprehensive Income Account.

Amount of total comprehensive income before considering gains/losses on investments = $164,972

Amount of total comprehensive income after considering gains/losses on investments = 164972 + 2592 + 13331

= $180,894.

Hence, the answer is B. $180894.